# Learning to engineer protein flexibility

### v1

@@ 4.2 ProteinMPNN can be steered by flexibility instructions
where $\bm{\hat{s}_{j}}$ is the predicted sequence for protein $j$ with backbone $\bm{\chi_{j}}$, $i\in\mathcal{M}_{j}$ indexes the mutated residues in the engineered region $S_{j}$, and $f^{native}_{ij}$ is the native flexibility of the residue (see Equation 1). The flexibility enrichment ratios $r_{ij}$ are then used to report the median enrichment ratio and the proportion of flexibility-increasing mutations, i.e., those for which $r_{ij}>1$.
@@ Results.
Table 4: Comparing the ability of flexibility engineering models to increase flexibility of engineered segments. The presented median enrichment ratios Med($r_{ij}$) and the proportion of flexibility-increasing mutations were obtained using the CATH4.3 test set with engineered segments of length $|S|=50$, and flexibility engineering instructions corresponding to native flexibility incremented by $\tau=5$ in the engineered segment. The best results in each category are highlighted in bold, second best are underlined.
@@
Interestingly, the most contributing component of Flexpert-Design is the inclusion of flexibility input and training with the Flexpert-3D generated flexibility pseudolabels. The additional optimization using $\mathcal{L}_{Flex}$ brings only relatively modest improvements in steerability (see Table 4 and Figure 8). Please compare “Flexpert-Design (without loss)”, which does not use $\mathcal{L}_{Flex}$ and the full model denoted “Flexpert-Design”. The results also clearly demonstrate significant improvements of our model in the ability to engineer flexibility compared to the ProteinMPNN baseline. Additional results in Section H of the Appendix also demonstrate virtually no loss in sequence recovery metrics of our approach compared to the ProteinMPNN baseline. The analysis of the structure preservation of the sequences with engineered increased flexibility is given in Appendix I.
@@ 5 Conclusions
In this work we address the problem of engineering protein flexibility—a highly relevant task due to its potential impact on the field of protein design. Firstly, we analyze the available data obtained in wet-lab experiments, physics-based simulations, or as a byproduct of machine learning-based structure prediction models. We identify the flexibility estimate by molecular dynamics simulation as the best learning target, set an upper bound on the potential performance of the learned predictor, and analyze the effect of the type of input protein structure on the performance of Anisotropic Network Models. Secondly, we develop a protein language model-based predictor of protein flexibility working either with protein sequence (Flexpert-Seq) or with both sequence and protein backbone structure on the input (Flexpert-3D) and show that our predictors outperform the relevant baselines. Thirdly, we propose a learning strategy Flexpert-Design, which uses the Flexpert-3D flexibility predictor to steer an inverse folding model toward a generation of sequences with increased flexibility. We demonstrate that inverse folding models can be steered toward generating protein sequences with increased flexibility as measured by our protein flexibility predictor. This opens up new possibilities for development of proteins with enhanced biological activities. In some practical protein engineering tasks, it would also be of interest to engineer protein sequences with decreased flexibility, where the ability of our current Flexpert-Design model seems limited (see Appendix K). This motivates future development of new models, which would provide a tighter control of the designed sequence’s flexibility and which would be successful regardless of whether the flexibility was instructed to increase or decrease. An interesting future direction could be to use Flexpert-3D or Flexpert-Seq for guidance of discrete flow matching models.
@@
The code together with the instructions how to download the data and trained weights for the reproduction of this work will be released at https://github.com/KoubaPetr/Flexpert.
@@ Appendix B Root mean square fluctuations (RMSF) as a measure of flexibility
The Root Mean Square Fluctuations (RMSF) are a way of quantifying protein flexibility, typically based on simulated data of dynamics or based on simplified models of dynamics such as the Elastic Network Models (ENM). In this work, we use RMSF computed on top of both MD simulations and ENMs (Anisotropic Network Models and Gaussian Network Models). These RMSFs were obtained in the following way:
@@ Appendix F Evaluation of Flexpert-3D and Flexpert-Seq on mdCATH dataset
Despite some performance drop with respect to the performance achieved on the ATLAS dataset (0.82 for Flexpert-3D and 0.76 for Flexpert-Seq), the performance of our predictors remains reasonable for the simulations performed at the temperature T = 320 K (0.66 for Flexpert-3D and 0.64 for Flexpert-Seq, for the topology filtered dataset). Considering that the temperature in this part of the mdCATH dataset is 20K away from the temperature of the simulations in the ATLAS training set, this appears as a reasonable generalization. We can notice that Flexpert-3D takes larger damage in its performance from the increased temperature of mdCATH simulations than Flexpert-Seq, suggesting that the sole reliance on sequence information makes the method more robust to the change in the simulation temperature. Note that the proteins are possibly getting unfolded in the high-temperature simulations of mdCATH, rendering non-physiological degrees of flexibility.
@@ Appendix G Qualitative analysis of Flexpert-Seq and Flexpert-3D flexibility predictions
In this section, we present some qualitative examples of Flexpert-Seq and Flexpert-3D predictions in the form of visualization of protein structures colored by the flexibility. The visualized flexibility is predicted by Flexpert-Seq and Flexpert-3D or estimated by MD (considered the ground truth) or by Anisotropic Network Models (ANM). See Figures 4 to 7.
@@ Appendix H Balancing sequence recovery and flexibility awareness in Flexpert-Design training
We tried to train Flexpert-Design from scratch using the $\mathcal{L}_{Flexpert}$ loss, but it seemed that the $\mathcal{L}_{Flex}$ loss was not informative in the early stages. Therefore, we decided to first train ProteinMPNN in a standard way for 100 epochs on topology split CATH4.3 dataset. Second, we finetuned the model for 12 hours on 1 GPU using the $\mathcal{L}_{Flexpert}$, which resulted in additional 8-9 epochs.
@@
Table 10 presents also the variant of Flexpert-Design trained using L1 loss instead of MSE, which seems to work better for the case of decreasing flexibility. However, despite achieving a modest decrease in flexibility, we consider this task as a limitation of Flexpert-Design. Our explanation is that Flexpert-Design is biased toward positive flexibility engineering instructions because it is trained on the native distribution of flexibilities, which are mostly positive as they approximate the RMSF quantity, which is by definition non-negative. Similarly, we suspect that the Flexpert-3D predictor might be slightly biased toward overpredicting flexibility since it predicts a flexibility increase even for the vanilla ProteinMPNN model. The possible reason is the effect of outliers in the training set. Outliers can exist in the dataset for positive values, but cannot exist for negative due to the non-negative nature of RMSF. Such disbalance can potentially skew the learning distribution toward higher values.

### v2

@@ 5.2 ProteinMPNN can be steered by flexibility instructions
where $\bm{\hat{s}_{j}}$ is the predicted sequence for protein $j$ with backbone $\bm{\chi_{j}}$, $i\in\mathcal{M}_{j}$ indexes the mutated residues in the engineered region $S_{j}$, and $f^{native}_{ij}$ is the native flexibility of the residue (see Equation (1)). The flexibility enrichment ratios $r_{ij}$ are then used to report the median enrichment ratio and the proportion of flexibility-increasing mutations, i.e., those for which $r_{ij}>1$.
@@ Results.
Table 4: Comparison of the ability of flexibility engineering models to increase flexibility of engineered segments. The presented median enrichment ratios Med($r_{ij}$) and the proportion of flexibility-increasing mutations were obtained using the CATH4.3 test set with engineered segments of length $|S|=50$, and flexibility engineering instructions corresponding to native flexibility incremented by $\tau=5$ in the engineered segment. The best results in each category are highlighted in bold, second best are underlined.
@@
Interestingly, the most contributing component of Flexpert-Design is the inclusion of flexibility input and training with the Flexpert-3D generated flexibility pseudolabels. The additional optimization using $\mathcal{L}_{Flex}$ brings only relatively modest improvements in steerability (see Table 4 and Figure 8). Please compare “Flexpert-Design (without loss)”, which does not use $\mathcal{L}_{Flex}$, and the full model denoted “Flexpert-Design”. The results also clearly demonstrate significant improvements of our model in the ability to engineer flexibility compared to the ProteinMPNN baseline. Additional results in Section H of the Appendix also demonstrate virtually no loss in sequence recovery metrics of our approach compared to the ProteinMPNN baseline. The analysis of the structure preservation of the sequences with engineered increased flexibility is given in Appendix I.
@@ 6 Conclusions
In this work, we address the problem of engineering protein flexibility—a highly relevant task due to its potential impact on the field of protein design. Firstly, we analyze the available data obtained in wet-lab experiments, physics-based simulations, or as a byproduct of machine learning-based structure prediction models. We identify the flexibility estimate by molecular dynamics simulation as the best learning target, set an upper bound on the potential performance of the learned predictor, and analyze the effect of the type of input protein structure on the performance of Anisotropic Network Models. Secondly, we develop a protein language model-based predictor of protein flexibility working either with protein sequence (Flexpert-Seq) or with both sequence and protein backbone structure on the input (Flexpert-3D) and show that our predictors outperform the relevant baselines. Thirdly, we propose a learning strategy Flexpert-Design, which uses the Flexpert-3D flexibility predictor to steer an inverse folding model toward a generation of sequences with increased flexibility. We demonstrate that inverse folding models can be steered toward generating protein sequences with increased flexibility as measured by our protein flexibility predictor. This opens up new possibilities for development of proteins with enhanced biological activities. In some practical protein engineering tasks, it would also be of interest to engineer protein sequences with decreased flexibility, where the ability of our current Flexpert-Design model seems limited (see Appendix K). This motivates future development of new models, which would provide a tighter control of the designed sequence’s flexibility and which would be successful regardless of whether the flexibility was instructed to increase or decrease. An interesting future direction could be to use Flexpert-3D or Flexpert-Seq for guidance of discrete flow matching models.
@@
The code, together with the instructions on how to download the data and trained weights for the reproduction of this work, can be found at https://github.com/KoubaPetr/Flexpert.
Acknowledgements
This work was supported by the Ministry of Education, Youth and Sports of the Czech Republic through projects ELIXIR (No. LM2023055), ESFRI RECETOX RI (No. LM2023069), and OP JAK CLARA (No. 02_23_029). We acknowledge VSB – Technical University of Ostrava, IT4Innovations National Supercomputing Center, Czech Republic, for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through the Ministry of Education, Youth and Sports of the Czech Republic through the e-INFRA CZ (No. 90254). This work was also supported by the European Union (ERC project FRONTIER (No. 101097822), ELIAS (No. 101120237), CLARA (No. 101136607)), the Technology Agency of the Czech Republic under the NCC Programme from the state budget (No. RETEMED TN02000122) and under the NRP from the EU RRF (No. TEREP TN02000122/001N), and the CETOCOEN EXCELLENCE (No. 857560) projects supported from the European Union’s Horizon 2020 research and innovation programme. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. Petr Kouba is a holder of the Brno Ph.D. Talent scholarship funded by the Brno City Municipality and the JCMM.
@@ Appendix B Root mean square fluctuations (RMSF) as a measure of flexibility
The Root Mean Square Fluctuations (RMSF) are a way of quantifying protein flexibility, typically based on simulated data of dynamics or based on simplified models of dynamics such as the Elastic Network Models (ENM). In this work, we use RMSF computed on top of both MD simulations and ENM (Anisotropic Network Models and Gaussian Network Models). These RMSFs were obtained in the following way:
@@ Appendix F Evaluation of Flexpert-3D and Flexpert-Seq on mdCATH dataset
Despite some performance drop with respect to the performance achieved on the ATLAS dataset (0.82 for Flexpert-3D and 0.76 for Flexpert-Seq), the performance of our predictors remains reasonable for the simulations performed at the temperature T = 320 K (0.66 for Flexpert-3D and 0.64 for Flexpert-Seq, for the topology filtered dataset). Considering that the temperature in this part of the mdCATH dataset is 20 K away from the temperature of the simulations in the ATLAS training set, this appears as a reasonable generalization. We also notice that Flexpert-3D experiences a more dramatic decrease in performance than Flexpert-Seq for higher temperatures in mdCATH simulations, suggesting that the sole reliance on sequence information makes the method more robust to the change in the simulation temperature. In addition, the proteins are possibly getting unfolded in the high-temperature simulations of mdCATH, leading to non-physiological degrees of flexibility.
@@ Appendix G Qualitative analysis of Flexpert-Seq and Flexpert-3D flexibility predictions
In this section, we present some qualitative examples of Flexpert-Seq and Flexpert-3D predictions in the form of visualization of protein structures colored by the flexibility. The visualized flexibility is predicted by Flexpert-Seq and Flexpert-3D or estimated by MD (considered as the ground truth) or by Anisotropic Network Models (ANM). See Figures 4 to 7.
@@ Appendix H Balancing sequence recovery and flexibility awareness in Flexpert-Design training
We tried to train Flexpert-Design from scratch using the $\mathcal{L}_{Flexpert}$ loss, but it seemed that the $\mathcal{L}_{Flex}$ loss was not informative in the early stages. Therefore, we decided to first train ProteinMPNN in a standard way for 100 epochs on topology split CATH4.3 dataset. Second, we finetuned the model for 12 hours on 1 GPU using the $\mathcal{L}_{Flexpert}$, which resulted in additional 8-11 epochs.
@@
Table 10 also presents the variant of Flexpert-Design trained using L1 loss instead of MSE, which seems to work better for the case of decreasing flexibility. However, despite achieving a modest decrease in flexibility, we consider this task as a limitation of Flexpert-Design. Our explanation is that Flexpert-Design is biased toward positive flexibility engineering instructions because it is trained on the native distribution of flexibilities, which are mostly positive as they approximate the RMSF quantity, which is by definition non-negative. Similarly, we suspect that the Flexpert-3D predictor might be slightly biased toward overpredicting flexibility since it predicts a flexibility increase even for the vanilla ProteinMPNN model. The possible reason is the effect of outliers in the training set. Outliers can exist in the dataset for positive values, but cannot exist for negative due to the non-negative nature of RMSF. Such disbalance can potentially skew the learning distribution toward higher values.